The functions are: f1(a,b,c,d) = ∑m(0,1,3,6,9,11) , f2(a,b,c,d) = Πm(2,3,4,5,7,10,11,12,13,14,15) and f3(a,b,c,d) = ab’c’ + a’b’c +b’cd

Complete the following problems for functions f1, f2, and f3

1. Implement the 3 functions on one(1) 4x16 decoder
Answer #1

f, Cabad) = Em (0,1,3,6,9,11) fy Ca.b,c,d) = Tim (2,3,4,5,7, 10, 11, 12,13,14,15) = {m(0,1, 6, 8, 9)fz (a, b, c,d) = abctäbct locd = abcd tabocłt ābcd tabcă. tabcd + a 5 cd = Em (2,3,8,9,11) . 4x16 obuo decoder - an slootoooo